# Capsule Network Projectors are Equivariant and Invariant Learners

### v3

@@ 4.1 Training Protocol
To directly compare with prior works employing the 3DIEBench dataset, we follow an identical training protocol, as defined in (Garrido et al., 2023). All methods employ a ResNet-18 encoder network ($f_{\theta}$), for the projection head ($h_{\phi}$) we compare a variety of hyperparameterisations, which we later describe in the following sections. For primary bench-marking we train our model for 2000 epochs using the Adam (Kingma & Ba, 2014) optimiser with default settings, a fixed learning rate of 1e-3 and batch size of 1024. For ablations and sensitivity analyses we train for 500 epochs and employ a batch size of 512, with other settings remaining unchanged. We find in practice that 500 epochs presents a strong correlation of performance. For all evaluations, pre-training was done with the equivariant criterion optimising for viewpoint rotation transformations, full details on these transformation groups, and the criterion are given in prior sections. By default the objective function weighting are as follows, $\lambda_{\text{inv}}=0.1,\;\lambda_{\text{equi}}=5,\;\lambda_{V}=10,\;\lambda_{C}=1$. Each self-supervised 2000 epoch pretraining run took approximately 22 hours using 3 Nvidia A100 80GB GPU’s, with 64 capsule models taking approximately 25 hours using 6 Nvidia A100 80GB GPU’s. All eval tasks are completed on a single Nvidia A100 80GB GPU and take approximately 6 hours for angle prediction, and 3 hours for classification.
Table 1: Evaluation of invariant properties via downstream classification task. Representations are learnt under the invariance and rotation equivariant objective, we evaluate both the representations and the intermediate embeddings of the projection head under varying number of capsules. FLOPs and Parameters correspond to computation during training, ‘-’ refers to non-compatible experiments. For non-capsule models, ‘All’, ‘Inv.’, and ‘Equi.’ refer to the full vector representation, the "left" half representation optimised under the invariant objective, and the "right" half representation optimised under the equivariant objective respectively. For capsule networks, ‘Inv.’, and ‘Equi.’ refer to the activation vector and pose matrices respectively.
@@ 4.2 Downstream Evaluation
To evaluate the quality of representations learnt under the invariant and viewpoint rotation equivariant self-supervised criterion, we use the standard benchmark approach of learning downstream task specific networks with frozen representations as input. In our case we evaluate the representations in three distinct tasks to evaluate both invariant and equivariant properties, we use a linear evaluation training protocol of the frozen representations. Further details of the evaluation protocol are given in the appendix.B.2.
Invariant Evaluation. To evaluate invariant properties of the representation we train a classifier on either the frozen representations output from the encoder network or the intermediate embeddings of the capsule network projector. Given our advocation for CapsNets we evaluate using both the standard linear classification and a capsule layer whose number of out capsules is set to the number of classes. All methods are trained for 300 epochs by cross entropy.
Equivariant Evaluation. Evaluating equivariant properties is achieved through a rotation prediction task in which a three layer MLP is trained to predict the quaternions defining the rotation between two views of the same object. We train for 300 epochs using MSE loss. Similarly to rotation prediction, we evaluate the representation’s equivariant properties by regressing the the colour hue of an object view. We train a single linear layer for 50 epochs using MSE. For evaluating the performance of predicting equivariance, we use the metric $R^{2}=1-\frac{\sum_{i}(y_{i}-\tilde{y}_{i})^{2}}{\sum_{i}(y_{i}-\bar{y})^{2}}$ where $\{y_{i}\}$ are the ground truth, target values of the equivariance, $\bar{y}$ is the mean value of these targets, and $\{\tilde{y}_{i}\}$ are the predictions. Higher $R^{2}$ indicates the model has a better fit for predicting equivariant transformations.
Table 2: Evaluation of equivariant properties via downstream rotation prediction (left) and colour prediction (right) tasks. Representations are learnt under the invariance and rotation equivariant objective, we evaluate both the representations and the intermediate embeddings of the projection head under varying number of capsules. ‘-’ refers to non-compatible experiments. For non-capsule models, ‘All’, ‘Inv.’, and ‘Equi.’ refer to the full vector representation, the "left" half representation optimised under the invariant objective, and the "right" half representation optimised under the equivariant objective respectively. For capsule networks, ‘Inv.’, and ‘Equi.’ refer to the activation vector and pose matrices respectively.
@@
Representation Quality. The performance of CapsIE for both invariant and equivariant benchmark tasks is given in Tables 1 and 2 respectively. We evaluate both the representations produced by the ResNet-18 encoder and the intermediate embeddings of the capsule layer projection head given different values for the numbers of capsules. We observe that across all models that the invariant properties captured within the representations marginally suffer compared to the MLP projector of SIE. This observation is expected given the significantly reduced number of embeddings employed in the invariant criterion compared to SIE. However, the evaluation of equivariant properties captured by the representations demonstrates that the use of a capsule projector in place of a MLP can lead to vastly improved performance in rotation prediction ($\uparrow 0.01\;R^{2}$) advancing the prior state-of-the-art whist also approaching the supervised baseline.
@@
Intermediate Projector Embeddings. The role of the projector is primarily employed to decorrelate the embeddings on which the objective function operates on from the representations employed downstream. The premise is to avoid representations that are over-fit to the self-supervised objective (Bordes et al., 2022). However, it has been well studied that it can be beneficial to maintain a number of projector layers and instead utilise intermediate projector embeddings for downstream tasks. Specifically, the equivariant information that has shown to be captured in the object pose (Ribeiro et al., 2022).
@@
We evaluate the intermediate embeddings output from the primary capsule layer in the same manner as the representations, however the activations and pose are given over a spatial region we perform average pooling to return an activation vector and a $4\times 4$ pose for each capsule which we then flatten into a vector. As with the representation evaluation we report our invariant and equivariant task performance in Tables 1 and 2 respectively. We find across all settings that evaluating the intermediate embeddings of the capsule projector leads to improved performance on rotation equivariant tasks compared to SIE intermediate embeddings. In our case, we demonstrate that the preservation of capsule layers for downstream tasks results in significantly closer evaluation performance to the representations, suggesting that capsule nets to presever such properties of interest compared to MLPs. Further investigations into this phenomena are left to future research.
@@ 4.3 Quantitative Evaluation of Equivariance
To quantitatively evaluate the equivariant performance of our method and capsule projector, we provide evaluations in line with those proposed in (Garrido et al., 2023). We report the Mean Reciprocal Rank (MRR) and Hit Rate at k (H@k) on the multi-object setting. Given a source and target pose of an object, we first compute the embeddings of each image, and pass the source embedding through the predictor and use the resulting vector to retrieve the nearest neighbours.
@@
All the results evaluated by the aforementioned metrics are given in Table 3. Our CapsIE network outperforms EquiMod, Only Equivariance and SIE by a considerable margin across all metrics and for all dataset splits. We achieve strong perfromance on PRE, reporting 0.21 PRE on the validation set compared to 0.48 for EquiMod and Only Equivariance and 0.29 for SIE. The same large gains in equivariant performance are shown for MRR and H@1 and H@5. Note, a random H@1 results in a performance for 2% (0.02) demonstrating that our method lies well above random.
Table 3: Quantitative evaluation of the predictor when using a Capsule network projector, using PRE, MRR and H@k. The source dataset which embeddings are computed and the dataset used for retrieval are given in the format source-retrieval for PRE and source for MRR and H@k. Here source refers to the set from which embeddings are computed (train or val), while retrieval corresponds to the set used for comparison/retrieval (train, val, or all = train + val).
@@ 4.4 Number of Capsules
Each capsule, in theory, should represent a unique concept, thus when the number of capsules is increased, logically so should the networks representation ability to capture an increasing number of semantic concepts. Observing the invariant performance during training (Figure 4) and downstream evaluation in Table 1, CapsIE gains slight improvement with the addition of more capsules. This demonstrates that our model has better utilised the additional representational power to improve performance. This pattern is also observed when evaluating equivariant properties (shown in Figure 4), yet is less pronounced. However, in Table 1 we also show that increasing the number of capsules of a supervised SR-Caps model trained in a standard supervised fashion is not an indicator of increased performance, aligning with prior capsule research (Everett et al., 2023). This differentiation in behaviour provides an interesting direction for future research.
@@
Figure 4: Invariant and equivariant performance of encoder representations during training when varying number of capsules. (left) Classification evaluation performance (top-1 %) and (right) Rotation prediction online evaluation performance ($R^{2}$) both learned encoder representations. This evaluation is co-optimised during training and hence differs from the downstream evaluation reported in the Table 1 and 2. We observe that as capsule numbers increase so does performance, however at low capsule numbers later in training invariant objective takes precedence.
@@ 5.1 Equivariant Self-Supervised Learning
Self-supervised learning has seen the majority of its success in the invariant setting by either contrastive (Chen et al., 2020), information maximisation (Zbontar et al., 2021; Bardes et al., 2021), or clustering methods (Caron et al., 2021; Assran et al., 2022). All families of approaches rely on training a network to be invariant to transformations by increasing the similarity between embeddings of the same image under augmentation. The differing approaches emerge from alternative methods to avoid collapse, the phenomena where embeddings fall into a lower-dimensional subspace rather than the entire available embedding space resulting in a trivial solution (Hua et al., 2021). Although these methods differ, they all produce similarly performing representations, hence we employ information maximisation methods as the basis of this work due to their computational efficiency.
Learning to be invariant to transformations is typically useful for semantic discrimination tasks, yet preserving information about the transformations can be highly beneficial. Some approaches have attempted to capture specific information regarding transformations by predicting the applied augmentation parameters (Lee et al., 2021), preserving the strength of augmentations (Xie et al., 2022) and introducing rotational transformations (Dangovski et al., 2022). However, as stated in (Garrido et al., 2023), these methods provide no guarantee that a mapping is learnt in the latent space that reflects the transformations in the input space. Hence, methods have been employed that address this limitation, (Devillers & Lefort, 2022; Park et al., 2022; Garrido et al., 2023) all employ predictor networks to predict the displacement representations in the latent space given one view representation and the transformation parameters. The latter, SIE (Garrido et al., 2023), is the basis of our work, which further extends prior methods by splitting representation vectors into invariant and equivariant parts to better separate differing information.

### v4

@@ 4.1 Training Protocol
To directly compare with prior works employing the 3DIEBench dataset, we follow an identical training protocol, as defined in Garrido et al. (2023). All methods employ a ResNet-18 encoder network ($f_{\theta}$). For the projection head ($h_{\phi}$), we compare various hyperparameterisations, which we describe in the following sections. For primary benchmarking we train our model for 2000 epochs using the Adam Kingma & Ba (2014) optimiser with default settings, a fixed learning rate of 1e-3 and a batch size of 1024. For ablations and sensitivity analyses we train for 500 epochs and employ a batch size of 512, with other settings remaining unchanged. We have found in practice that 500 epochs presents a strong correlation with performance. For all evaluations, pre-training was done with the equivariant criterion optimising for viewpoint rotation transformations. Full details on these transformation groups and the criteria are given in prior sections. By default the objective function weighting are as follows, $\lambda_{\text{inv}}=0.1,\;\lambda_{\text{equi}}=5,\;\lambda_{V}=10,\;\lambda_{C}=1$. Each self-supervised 2000-epoch pretraining run took approximately 22 hours using three Nvidia A100 80GB GPUs for the 32 capule model, whereas the 64 capsule models, and required approximately 25 hours using six Nvidia A100 80GB GPUs. For comparison SIE training took approximately 26 hours using three Nvidia A100 80GB GPUs. All evaluation tasks are completed on a single Nvidia A100 80GB GPU and take approximately 6 hours for angle prediction, and 3 hours for classification. Given the computational overheads involved, all results are presented as a single run where the seed is set to “2224” with exception to Table. 4 where the mean and standard deviation of 5 seeds are reported.
Table 1: Evaluation of invariant properties via downstream classification task. Representations are learnt under the invariance and rotation equivariant objective. We evaluate both the representations and the intermediate embeddings of the projection head under varying numbers of capsules. FLOPs and Parameters correspond to computation during training, ‘-’ refers to non-compatible experiments. For non-capsule models, ‘All’, ‘Inv.’, and ‘Equi.’ refer to the full vector representation, the "left" half representation optimised under the invariant objective, and the "right" half representation optimised under the equivariant objective, respectively. For capsule networks, ‘Inv.’, and ‘Equi.’ refer to the activation vector and pose matrices, respectively.
@@ 4.2 Downstream Evaluation
To evaluate the quality of representations learnt under the invariant and viewpoint rotation equivariant self-supervised criterion, we use the standard benchmark approach of learning downstream task-specific networks with frozen representations as input. In our case, we evaluate the representations in three distinct tasks to evaluate both invariant and equivariant properties. We use a linear evaluation training protocol of the frozen representations. Further details of the evaluation protocol are given in the appendix A.3.2.
Invariant Evaluation. To evaluate invariant properties of the representation, we train a classifier on either the frozen representations output from the encoder network or the intermediate embeddings of the capsule network projector. Given our advocacy for CapsNets, we evaluate using both the standard linear classification and a capsule layer whose number of output capsules is set to the number of classes. All methods are trained for 300 epochs by cross entropy.
Equivariant Evaluation. Evaluating equivariant properties is achieved through a rotation prediction task in which a three layer MLP is trained to predict the quaternions defining the rotation between two views of the same object. We train for 300 epochs using MSE loss. Similar to rotation prediction, we evaluate the representation’s equivariant properties by regressing the colour hue of an object view. We train a single linear layer for 50 epochs using MSE. For evaluating the performance of predicting equivariance, we use the metric $R^{2}=1-\frac{\sum_{i}(y_{i}-\tilde{y}_{i})^{2}}{\sum_{i}(y_{i}-\bar{y})^{2}}$ where $\{y_{i}\}$ are the ground truth (target values of the equivariance), $\bar{y}$ is the mean value of these targets, and $\{\tilde{y}_{i}\}$ are the predictions. Higher $R^{2}$ indicates the model has a better fit for predicting equivariant transformations.
Table 2: Evaluation of equivariant properties via downstream rotation prediction (left) and colour prediction (right) tasks. Representations are learnt under the invariance and rotation equivariant objective, we evaluate both the representations and the intermediate embeddings of the projection head under varying number of capsules. ‘-’ refers to non-compatible experiments. For non-capsule models, ‘All’, ‘Inv.’, and ‘Equi.’ refer to the full vector representation, the "left" half representation optimised under the invariant objective, and the "right" half representation optimised under the equivariant objective, respectively. For capsule networks, ‘Inv.’, and ‘Equi.’ refer to the activation vector and pose matrices, respectively.
@@
Representation Quality. The performance of CapsIE for both invariant and equivariant benchmark tasks is given in Tables 1 and 2, respectively. We evaluate both the representations produced by the ResNet-18 encoder and the intermediate embeddings of the capsule layer projection head given different values for the number of capsules. We observe that across all models that the invariant properties captured within the representations marginally suffer compared to the MLP projector of SIE. This observation is expected given the significantly reduced number of embeddings employed in the invariant criterion compared to SIE. However, the evaluation of equivariant properties captured by the representations demonstrates that the use of a capsule projector in place of an MLP can lead to marginally improved performance in rotation prediction ($\uparrow 0.01\;R^{2}$) advancing the prior state-of-the-art whist also approaching the supervised baseline.
Additionally, we include a colour prediction task as a diagnostic test to determine whether the capsule projector additionally encodes colour changes, an additional group element, without being explicitly optimised to do so. We hypothesise that poor colour prediction from CapsIE’s pose is due to its optimisation being directly for angle and pose changes only. Since we observe these results with $R^{2}$ slightly below zero. However, given the negative score is very marginal, around the zero mark, we do not interpret this as a significant negative correlation. Future investigations into the more significant negative correlation of the 64-capsule model, and whether the score negatively correlates with the number of capsules, are warranted. In Table 2, we can conclude that capsules do not inherently encode colour information under such strong objectives. This is beneficial as it allows for more finegrained control for properties that are optimised for and does not arbritarily encode additional group elements.
@@
Intermediate Projector Embeddings. The role of the projector is primarily employed to decorrelate the embeddings on which the objective function operates on from the representations employed downstream. The premise is to avoid representations that are over-fit to the self-supervised objective (Bordes et al., 2023). However, it has been well studied that it can be beneficial to maintain a number of projector layers and instead utilise intermediate projector embeddings for downstream tasks. Specifically, the equivariant information that has been shown to be captured in the object pose (De Sousa Ribeiro et al., 2024).
@@
We evaluate the intermediate embeddings output from the primary capsule layer in the same manner as the representations; however, the activations and pose are given over a spatial region, so we perform average pooling to return an activation vector and a $4\times 4$ pose for each capsule, which we then flatten into a vector. As with the representation evaluation, we report our invariant and equivariant task performance in Tables 1 and 2, respectively. We find that evaluating the intermediate embeddings of the capsule projector consistently leads to improved performance on rotation-equivariant tasks compared to SIE intermediate embeddings across all settings. In our case, we demonstrate that the preservation of capsule layers for downstream tasks results in significantly better evaluation performance of the representations, suggesting that capsule nets preserve such properties of interest compared to MLPs. Further investigations into this phenomenon are left to future research.
@@ 4.3 Quantitative Evaluation of Equivariance
To quantitatively evaluate the equivariant performance of our method and capsule projector, we provide evaluations in line with those proposed in Garrido et al. (2023). We report the Mean Reciprocal Rank (MRR) and Hit Rate at k (H@k) on the multi-object setting. Given a source and target pose of an object, we first compute the embeddings of each image and pass the source embedding through the predictor. We then use the resulting vector to retrieve the nearest neighbours.
@@
All the results evaluated by the aforementioned metrics are given in Table 3. Our CapsIE network outperforms EquiMod, Only Equivariance and SIE by a considerable margin across all metrics and for all dataset splits. We achieve strong perfromance on PRE, reporting 0.21 PRE on the validation set compared to 0.48 for EquiMod and Only Equivariance and 0.29 for SIE. The same significant gains in equivariant performance are shown for MRR and H@1 and H@5. Note, a random H@1 results in a performance of 2% (0.02), demonstrating that our method lies well above random.
Table 3: Quantitative evaluation of the predictor when using a Capsule network projector, using PRE, MRR and H@k. The source dataset, for which embeddings are computed, and the dataset used for retrieval are given in the format source-retrieval for PRE and source for MRR and H@k. Here source refers to the set from which embeddings are computed (train or val), while retrieval corresponds to the set used for comparison/retrieval (train, val, or all = train + val).
@@ 4.4 Number of Capsules
Each capsule, in theory, should represent a unique concept; thus when the number of capsules is increased, logically so should the network’s representation ability to capture an increasing number of semantic concepts. Observing the invariant performance during training (Figure 4) and downstream evaluation in Table 1, CapsIE gains a slight improvement with the addition of more capsules. Here, online training refers to a co-optimisation scheme where an evaluation network is trained in parallel during standard pretraining procedure, this allows for analysis and visualisation of training dynamics rather than relying on the losses of each term directly. Since this evaluation network has been trained for far longer and on a much more diverse representation set during training, the performance is typically higher. The performance however, demonstrates that our model has better utilised the additional representational power to improve performance. This pattern is also observed when evaluating equivariant properties (shown in Figure 4), yet it is less pronounced. However, in Table 1 we also show that increasing the number of capsules in a supervised SR-Caps model trained in a standard supervised fashion is not an indicator of increased performance, aligning with prior capsule research (Everett et al., 2023). This differentiation in behaviour provides an interesting direction for future research.
@@
Figure 4: Invariant and equivariant performance of encoder representations during training when varying the number of capsules. (left) Classification evaluation performance (top-1 %) and (right) Rotation prediction online evaluation performance ($R^{2}$) both learned encoder representations. This evaluation is co-optimised during training and hence differs from the downstream evaluation reported in the Table 1 and 2. We observe that as capsule numbers increase so does performance; however, at low capsule numbers, later in training invariant objective takes precedence.
4.5 Additional Quantitative Results
4.5.1 Objaverse
Datasets such as 3DIEBench, derived from ShapeNet Chang et al. (2015), have some limitations, considering that they are synthetically generated, although photorealistic. However, datasets such as Objaverse (Deitke et al., 2023) contain real-world scans, therefore providing the possibility of evaluating equivariant models on different settings. To further validate CapsIE on a randomly rotated multi-view dataset, we used a subset of Objaverse-LVIS Deitke et al. (2023) with six classes (airplane, bench, car automobile, chair, coffee table, and gun), following the class selection of Wang et al. (2024). We evaluated classification and rotation prediction using two approaches. Firstly, we perform transfer learning by fine-tuning the entire network pre-trained on 3DIEBench as described previously. Secondly, we take the model pre-trained on 3DIEBench, freeze the backbone encoder, and train only the task-specific heads from scratch following the same evaluation procedure as described in A.3.2.
Results shown in Table 4 demonstrate that CapsIE performs competitively against other methods and, in some cases, outperforms them by a large margin, e.g. rotation with a frozen backbone. We do note that in the fine-tuning case our capsule architecture outperforms the SIE invariant baseline demonstrating that the learned invariant representations may be generally well expressed but need a further small amount of supervised tuning to better structure the representations more effectively. This introduces a promising direction for future work, and the applicability of pretrained capsule networks.
Table 4: Evaluation on a subset of Objaverse-LVIS using a ResNet-18 backbone. Representations are evaluated on an invariant task (classification) and an equivariant task (rotation prediction). Each model is evaluated five times with different random seeds.
Pre-training (Frozen Backbone) Transfer Learning (Fine-tuning)
Method Classification (Top-1) Rotation ($R^{2}$) Classification (Top-1) Rotation ($R^{2}$)
[…]
4.5.2 MOVi-E
To assess performance under more realistic and challenging conditions, we evaluate our approach on the Multi-Object Video (MOVi-E) dataset Greff et al. (2022). MOVi-E is a synthetic benchmark containing scenes with up to 17 distinct objects placed in photorealistic environments, often with occlusions and complex backgrounds. Each scene is generated through a 2-second rigid-body simulation in which multiple objects fall and interact. While MOVi-E provides sequences with a linearly moving camera, in our experiments, we sample individual frames and process them independently rather than using the temporal dimension. More information on the dataset can be found in the original paper Greff et al. (2022) and the public repository [^1] https://github.com/google-research/kubric/blob/main/challenges/movi.
Table 5: Transfer learning via DETR fine-tuning with frozen backbone on MOVi-E.
Method Classification (Top-1) Rotation ($R^{2}$) mAP mAP50 mAP75
SIE 73.7 0.20 26.47 41.83 28.26
CapsIE 72.5 0.23 28.26 43.41 31.38
In this setting, the task involves detecting every object (via bounding boxes), classifying its type, and estimating its pose relative to the camera frame. For object detection, we adopt the DETR architecture Carion et al. (2020), initialising the ResNet-50 backbone with weights pre-trained on 3DIEBench. The backbone is frozen, while the transformer encoder, decoder, and prediction heads are fine-tuned on MOVi-E. To enable pose estimation, we extend DETR with an additional MLP head that regresses rotation quaternions. This predictor mirrors the bounding box regression module, using a three-layer MLP with 256 hidden units. The quaternion regression loss is defined as mean squared error and added to the overall training objective, which is a weighted sum of individual DETR losses Carion et al. (2020).
Training is carried out for 200 epochs with a batch size of 64, starting from a learning rate of 0.0001 that is reduced tenfold after epoch 100. The quaternion loss is assigned a weight of 2, and the Generalised Intersection over Union loss a weight of 3, while all other hyperparameters remain unchanged from the original DETR setup.
The results of classification, detection, and rotation regression are summarised in Table  5. Our method outperforms SIE in all tasks except classification, and it is the strongest competing equivariant approach, indicating strong generalisation and robustness to complex multi-object scenarios. Although CapsIE performs competitively, its performance remains constrained by the difficulty of MOVi-E and the fact that the ResNet-50 backbone was trained only in single-object settings. We view this as a promising direction for future work, particularly through adapting the method to exploit temporal video information.
@@ 5.1 Equivariant Self-Supervised Learning
Self-supervised learning has seen the majority of its success in the invariant setting by contrastive (Chen et al., 2020), information maximisation (Zbontar et al., 2021; Bardes et al., 2022), or clustering methods (Caron et al., 2021; Assran et al., 2022). All families of approaches rely on training a network to be invariant to transformations by increasing the similarity between embeddings of the same image under augmentation. The differing approaches emerge from alternative methods to avoid collapse, a phenomenon where embeddings fall into a lower-dimensional subspace rather than the entire available embedding space, resulting in a trivial solution (Hua et al., 2021). Although these methods differ, they all produce similarly performing representations, hence we employ information maximisation methods as the basis of this work due to their computational efficiency.
Learning to be invariant to transformations is typically useful for semantic discrimination tasks, yet preserving information about the transformations can be highly beneficial. Some approaches have attempted to capture specific information regarding transformations by predicting the applied augmentation parameters (Lee et al., 2021), preserving the strength of augmentations (Xie et al., 2022) and introducing rotational transformations (Dangovski et al., 2022). However, as stated in Garrido et al. (2023), these methods provide no guarantee that a mapping is learnt in the latent space that reflects the transformations in the input space. Hence, methods have been employed that address this limitation (Devillers & Lefort, 2023; Park et al., 2022; Garrido et al., 2023). All of these methods employ predictor networks to predict displacement representations in the latent space, given a single view representation and the transformation parameters. The latter, SIE Garrido et al. (2023), is the basis of our work, which further extends prior methods by splitting representation vectors into invariant and equivariant parts to better separate differing information.